OpenStudy (anonymous):

FIND THE EQUATION FOR THE LINE THROUGH (2,5) AND (1,9)

OpenStudy (anonymous):

y=-4x+13

jimthompson5910 (jim_thompson5910):

m=(y2-y1)/(x2-x1) m=(9-5)/(1-2) m=(4)/(1-2) m=(4)/(-1) m=-4 So the slope of the line that goes through the points (2,5) and (1,9) is m=-4 Now let's use the point slope formula: y-y1=m(x-x1) y-5=-4(x-2) y-5=-4x+-4(-2) y-5=-4x+8 y=-4x+8+5 y=-4x+13 So the answer is y = -4x+13